a wandering grizzly bear travels 563 meters due west, then meanders a distance of 358 meters in direction 32 degrees south of west, followed by a 208 meters distance directly south. find the magnitude of the distance that would be needed for the grizzly bear return straight to its original starting point? find the angle relative to east that the grizzly would need to head in order to return straight to its original starting point?
1
Expert's answer
2020-12-21T10:43:52-0500
"s_x=563+358\cos{32}=866.6\ m\\s_y=208+358\sin{32}=397.7\ m\\s=\sqrt{866.6^2+397.7^2}=953\ m"

"\theta=\arctan{\frac{397.7}{866.6}}=24.65\degree \text{ N of E}"